A Carnot heat engine is operated between two heat reservoirs at temperature 400K and 300K. If the engine receives 200 Caloriesfrom the 400K reservoir, what is the thermal efficiency of the engine
1
Expert's answer
2020-12-06T17:27:25-0500

According to the Carnot's theorem, the thermal efficiency efficiency of a Carnot heat engine is the ratio of the temperature difference between the reservoirs and the temperature of the hot reservoir, expressed in the equation (https://en.wikipedia.org/wiki/Carnot%27s_theorem_(thermodynamics)):



"\eta =\dfrac{T_2-T_1}{T_2} = 1-\dfrac{T_1}{T_2}"


where, in our case, "T_1 = 300K" and "T_2 = 400K". Thus, obtain:



"\eta = 1-\dfrac{300K}{400K} = 0.25"

Answer. 0.25.